### v2

# Almost equivalences between Tamarkin category and Novikov sheaves

Abstract
We revisit the relationship between the Tamarkin’s extra variable $\mathbb{R}_{t}$ and Novikov rings. We prove that the equivariant version of Tamarkin category is almost equivalent (in the sense of almost mathematics) to the category of derived complete modules over the Novikov ring.
1 Introduction
In symplectic geometry, there exists a series of extra variables, which are (expected to be) just various aspects of one variable.
  1. 1.
    In microlocal sheaf theory, it is called $t$ and introduced by Tamarkin [Tam18], which was also envsioned by Sato in relation with WKB analysis [SK14].
  2. 2.
    In deformation quantization/WKB analysis/twistor theory, it is the Laplace dual of the inverse of $\hbar$.
  3. 3.
    In Floer theory, it is the exponent/valuation of the universal Novikov ring.
For example, in [Kuwal], we discussed the relationship between (1) and (2). Also, the relationship between (1) and (3) can be seen from the work of Tamarkin, and later clarified by [AI, IK].
In this paper, we investigate the relationship between (1) and (3) further.
In [Kuwal], we found the Novikov ring a posteriori after the construction of the equivariant version of Tamarkin’s category. In this paper, we try to build the Novikov ring a priori into sheaves. Hence, it gives an alternative model of the category of sheaf quantizations.
We would like to formulate our main theorem now. Let $\mathbb{K}$ be a field. Let $M$ be a manifold and $\mathbb{R}_{t}$ be the real line. We denote the discrete additive group of the real numbers by $\mathbb{R}_{d}$. We then have the equivariant derived category $\mathrm{Sh}^{\mathbb{R}_{d}}(M\times\mathbb{R}_{t},\mathbb{K})$ of $\mathbb{K}$-modules sheaves. We quotient this category by the sheaves with non-positive microsupport and denote it by $\mu(T^{*}M,\mathbb{K})$. On the other hand, we have the universal Novikov ring $\Lambda_{0}$ over $\mathbb{K}$. We denote the derived category of $\Lambda_{0}$-module sheaves by $\mathrm{Sh}(M,\Lambda_{0})$.
Theorem 1.1.
We have an almost embedding
$\mathfrak{A}\colon\mu(T^{*}M)\hookrightarrow_{\mathrm{a}}\mathrm{Sh}(M,\Lambda_{0}).$ (1.1)
The almost image satisfies the derived completeness.
Here the term “almost” comes from the almost mathematics [GR03]. Roughly speaking, the above theorem without “almost” holds after negating almost zero modules. The precise meaning will be explained in the body of the paper. The use of the almost mathematics in symplectic geometry also happened in [Fuk21].
In the body of paper, we will state a more generalized version of the theorem: Namely, for sheaves equivariant with respect to a subgroup $\mathbb{G}\subset\mathbb{R}$. In this version, the right hand side is a sheaf valued in modules over a certain completion of infinite versions of $A_{n}$-quiver algebra. For example, in the classical case, it is a certain completion of the poset $(\mathbb{R},\geq)$ used in the literature of persistent modules.
A variant of the main theorem was also described in [Kuw21] where we used the finite Novikov ring $\mathbb{K}[\mathbb{R}_{\geq 0}]$, and compared it with the enhanced ind-sheaves, to reformulate the holonomic Riemann–Hilbert correspondence [DK16].
When we try to relate (1) and (3) categorically, it is easier to construct a functor valued in the sheaves of modules over the Novikov ring. For example, our construction is crucial when we would like to relate sheaf theory and Fukaya category. In the subsequent publication, we plan to combine this result with a generalization of Viterbo’s construction [Vit] to upgrade the Fukaya–sheaf correspondence [GPS20] to an equivalence over the Novikov ring. An integral version of the statement is carried out in a joint work with Petr and Shende [KPS] in a different way, but we use some computations from this article. Also, we plan to rewrite the preprint [Kuw22] using the present formalism.
Aside from the above applications, in this paper,we will explain two applications of (the philosophy) of the theorem.
The first one is a first step toward nonconic microlocal sheaf theory of sheaves valued in modules over real valuation rings. Namely, for a manifold $M$, a real valuation ring $R$, and a sheaf $\mathcal{E}$ of modules over $R$, we can define a subset $\mu\mathrm{supp}(\mathcal{E})\subset T^{*}M$ by interpreting Tamarkin’s non-conic microsupport. This refines Kashiwara–Schapira’s microsupport [KS94].
The second one is an introduction of curved sheaves and twisted sheaves. When we would like to relate sheaves with Floer theory, such notions are very natural, since Floer complex can be curved and allowed to be deformed by some bulk classes.
Notation
Let $\mathbb{K}$ be a field.
Acknowledgment
This work is supported by JSPS KAKENHI Grant Numbers 22K13912, 23H01068, and 20H01794. I’d like to thank Yuichi Ike for related discussions. I also thank Tomohiro Asano for pointing out a mistake.
2 Novikov rings
In this section, we define Novikov rings and explain some properties.
2.1 Definition
Let $\mathbb{R}$ be the 1-dimensional Euclidean vector space. Let $\mathbb{G}$ be a subgroup of $\mathbb{R}$. Then $\mathbb{R}_{\geq 0}\cap\mathbb{G}$ has a semigroup structure with respect to the addition. We denote the corresponding polynomial ring by $\mathbb{K}[\mathbb{R}_{\geq 0}\cap\mathbb{G}]$. We denote the indeterminate corresponding to $a\in\mathbb{R}_{\geq 0}\cap\mathbb{G}$ by $T^{a}$. Let ||$\cdot$|| ⋅ | be the Euclidean norm of $\mathbb{R}$. For $r\in\mathbb{R}_{>0}$, we denote the ideal of $\mathbb{K}[\mathbb{R}_{\geq 0}\cap\mathbb{G}]$ generated by $T^{a}$’s with $a>r$ by $\mathfrak{m}(r)$. Obviously, $\mathfrak{m}(r^{\prime})\supset\mathfrak{m}(r)$ if $r>r^{\prime}$. Hence $\mathbb{K}[\mathbb{R}_{\geq 0}\cap\mathbb{G}]/\mathfrak{m}(r)$ forms a projective system.
Definition 2.1.
The Novikov ring $\Lambda_{0}^{\mathbb{G}}$ associated to $\mathbb{G}$ is defined by
$\Lambda_{0}^{\mathbb{G}}:=\lim_{\begin{subarray}{c}\longleftarrow\\ r\rightarrow+\infty\end{subarray}}\mathbb{K}[\mathbb{R}_{\geq 0}\cap\mathbb{G}]/\mathfrak{m}(r).$ (2.1)
Example 2.2 (The universal Novikov ring).
If $\mathbb{G}=\mathbb{R}$, the definition can be read as follows: Consider the semigroup of the non-negative real numbers $\mathbb{R}_{\geq 0}$. We consider the polynomial ring $\mathbb{K}[\mathbb{R}_{\geq 0}]$. We denote the indeterminate corresponding to $a\in\mathbb{R}_{\geq 0}$ by $T^{a}$. We set
$\Lambda_{0}:=\Lambda_{0}^{\mathbb{R}}:=\lim_{\begin{subarray}{c}\longleftarrow\\ a\rightarrow+\infty\end{subarray}}\mathbb{K}[\mathbb{R}_{\geq 0}]/T^{a}\mathbb{K}[\mathbb{R}_{\geq 0}].$ (2.2)
The ring is very useful in symplectic topology [FOOO09] to control the energy/disk area.
Example 2.3.
Similarly, if $\mathbb{G}=\mathbb{Z}$, we get the formal power series ring $\Lambda_{0}^{\mathbb{Z}}\cong\mathbb{K}[[T]]$.
Example 2.4.
Similarly, if $\mathbb{G}=\mathbb{O}:=\{0\}$, we get $\Lambda_{0}^{\mathbb{O}}\cong\mathbb{K}$.
We list up some properties of $\Lambda_{0}^{\mathbb{G}}$.
Lemma 2.5.
For any $\mathbb{G}$, the ring $\Lambda_{0}^{\mathbb{G}}$ is
  1. 1.
    an integral domain, and
  2. 2.
    a local ring.
Proof.
These are obvious. The maximal ideal $\mathfrak{m}$ is given by the ideal generated by $\left\{T^{a}\mathrel{}\middle|\mathrel{}a\in\mathbb{G}\cap\mathbb{R}_{>0}\right\}$. ∎
2.2 Modules over the Novikov ring
Let $R$ be a Novikov ring i.e., $R=\Lambda_{0}^{\mathbb{G}}$ for some $\mathbb{G}$. We denote the abelian category of $R$-modules by $\mathrm{Mod}^{\heartsuit}(R)$. We also denote its derived category by $\mathrm{Mod}(R)$. We regard it as an $\infty$-category (or more concretely, a dg-category).
Let $M$ be a topological space. Let $R_{M}$ be the constant sheaf valued in $R$. We denote the abelian category $R_{M}$-modules by $\mathrm{Mod}^{\heartsuit}(R_{M})$. We similarly consider the derived category of $\mathrm{Mod}^{\heartsuit}(R_{M})$, and denote it by $\mathrm{Mod}(R_{M})$, viewed as an $\infty$-category (or more concretely, a dg-category).
In the case of $\mathbb{G}\neq\mathbb{R}$, to compare with sheaf theory, we introduce a noncommutative algebra as follows. We first introduce the following quiver: The set of vertices is $\mathbb{R}/\mathbb{G}$. For $[c],[c^{\prime}]\in\mathbb{R}/\mathbb{G}$, the set of arrows from $[c]$ to $[c^{\prime}]$ is identified with the set
$\left\{d\in\mathbb{R}_{\geq 0}\mathrel{}\middle|\mathrel{}[c+d]=[c^{\prime}]\right\}.$ (2.3)
We denote the morphism corresponding to $d$ by $e_{[c],d}$. We denote the associated quiver algebra by $\mathbb{K}[Q(\mathbb{R}/\mathbb{G})]$. An element of this algebra is of the form
$\prod_{[c]\in\mathbb{R}/\mathbb{G}}\sum_{d\in\mathbb{R}_{\geq 0}}a_{[c],d}e_{[c],d}$ (2.4)
where $a_{[c],d}\in\mathbb{K}$ are zero except for finitely many $d$ for each $[c]$. For two elements, $\prod_{[c]\in\mathbb{R}/\mathbb{G}}\sum_{d\in\mathbb{R}_{\geq 0}}a_{[c],d}e_{[c],d},\prod_{[c]\in\mathbb{R}/\mathbb{G}}\sum_{d\in\mathbb{R}_{\geq 0}}b_{[c],d}e_{[c],d}$, we define the product by
$\left(\prod_{[c]\in\mathbb{R}/\mathbb{G}}\sum_{d\in\mathbb{R}_{\geq 0}}b_{[c],d}e_{[c],d}\right)\cdot\left(\prod_{[c]\in\mathbb{R}/\mathbb{G}}\sum_{d\in\mathbb{R}_{\geq 0}}a_{[c],d}e_{[c],d}\right)=\prod_{[c]\in\mathbb{R}/\mathbb{G}}\sum_{d^{\prime\prime}\in\mathbb{R}_{\geq 0}}\left(\sum_{d+d^{\prime}=d^{\prime\prime}}b_{[c+d],d^{\prime}}a_{[c],d}\right)e_{[c],d^{\prime\prime}}$ (2.5)
The sums are finite sums, so this is well-defined.
For each $c$, the identity morphism is denoted by $e_{c}\in\mathbb{K}[Q(\mathbb{R}/\mathbb{G})]$. For $\ell>0$, we denote the ideal generated by the arrows represented by the positive numbers greater than $\ell$ by $\mathfrak{m}(\ell)$. We take the completion of the algebra with $\mathfrak{m}(\ell)$-adic topology and denote it by $L_{0}^{\mathbb{G}}$. More explicitly, an element of $L_{0}^{\mathbb{G}}$ is of the form
$\prod_{[c]\in\mathbb{R}/\mathbb{G}}\sum_{d\in\mathbb{R}_{\geq 0}}a_{[c],d}e_{[c],d}$ (2.6)
Here, for each $[c]$ and $L>0$, $a_{[c],d}\in\mathbb{K}$ with $d\leq L$ are zero except for finitely many $d$ i.e., “Novikov sum”. The obvious multiplication is again well-defined.
By definition, the following is obvious:
Lemma 2.6.
As $\mathbb{K}$-modules, $L_{0}^{\mathbb{G}}\cong\prod_{c\in\mathbb{R}/\mathbb{G}}\Lambda_{0}.$
We now would like to see several examples.
Example 2.7.
The case of $\mathbb{R}=\mathbb{G}$. Since $\mathbb{R}/\mathbb{G}=\{*\}$, the quiver algebra is simply $\mathbb{K}[Q(\mathbb{R}/\mathbb{G})]=\mathbb{K}[\mathbb{R}_{\geq 0}]$. Hence the completion is $L_{0}^{\mathbb{G}}=\Lambda_{0}$.
Example 2.8.
Let $M$ be a finite-dimensional persistence module, namely, a functor from the poset category $(\mathbb{R},\geq)$ to the category of finite-dimensional vector spaces. For each $c\in\mathbb{R}$, we denote the image under the functor by $M_{c}$. We suppose that there exists $L\in\mathbb{R}$ such that $M_{c}=0$ for any $c<L$. In the following, we will see that $\prod_{c\in\mathbb{R}}M_{c}$ carries a $L_{0}^{\mathbb{O}}$-module structure.
For $c\leq c^{\prime}$, we denote the structure morphism by $t_{c,c^{\prime}}\colon M_{c}\rightarrow M_{c^{\prime}}$. We set $N_{-c}:=M_{c}$. Take an element $\prod_{c\in\mathbb{R}}n_{c}=\prod_{c>-L}n_{c}\in\prod_{c\in\mathbb{R}}N_{c}$. We also set $e_{[c],d}=:f_{c,c+d}$. The action is defined by
$\left(\prod_{c\in\mathbb{R}}\sum_{c\leq c^{\prime}}a_{c,c^{\prime}}f_{c,c^{\prime}}\right)\cdot\prod_{c\in\mathbb{R}}n_{c}=\prod_{c\in\mathbb{R}}n^{\prime}_{c}$ (2.7)
where
$M_{-c}:=N_{c}\ni n_{c}^{\prime}=\sum_{c\leq c^{\prime}}a_{c,c^{\prime}}t_{-c^{\prime},-c}(n_{c^{\prime}}).$ (2.8)
If $c^{\prime}$ is sufficiently large, $n_{c^{\prime}}=0$. Hence the sum (2.8) is a finite sum, hence well-defined.
The case of $\Lambda_{0}$ (i.e., $\mathbb{G}=\mathbb{R}$) is universal in the following sense:
Lemma 2.9.
We have an action of $\Lambda_{0}$ on $L_{0}^{\mathbb{G}}$. In particular, we have the forgetful functor
$\mathfrak{f}\colon\mathrm{Mod}(L_{0}^{\mathbb{G}})\rightarrow\mathrm{Mod}(\Lambda_{0}).$ (2.9)
Proof.
On $e_{[c],d}$, $T^{d^{\prime}}\in\Lambda_{0}$ acts on it by $e_{[c],d}\mapsto e_{[c],d+d^{\prime}}$. This defines the desired action. ∎
2.3 Real valuation
We will later use the notion of real valuation.
Definition 2.10.
Let $A$ be an integral domain with a map $v\colon A\rightarrow\mathbb{R}_{\geq 0}$. We say $(A,\mathfrak{v})$ is a real valuation ring if
  1. 1.
    $v(0)=0,v(-a)=v(a)$,
  2. 2.
    $v(a+b)\geq v(a)+v(b)$,
  3. 3.
    $v(ab)=v(a)v(b)$
for any $a,b\in A$.
For the Novikov ring $\Lambda_{0}$, we set
$v(x):=\min\left\{c\in\mathbb{R}_{\geq 0}\mathrel{}\middle|\mathrel{}a\in T^{c}\mathfrak{m}\right\}.$ (2.10)
This obviously gives a real valuation of $\Lambda_{0}$.
3 Derived complete modules
In this section, we recall some basic properties of derived complete modules. Our references are [Ked] and [Aut].
3.1 Derived completeness
We first recall the definition of the completeness/derived completeness.
Definition 3.1.
Let $A$ be a ring and $I$ be a finitely generated ideal of $A$. Let $M$ be an $A$-module. The inverse limit ${\lim_{\begin{subarray}{c}\longleftarrow\\ n\rightarrow\infty\end{subarray}}M/I^{n}M}$ is called the $I$-adic completion of $M$.
We say $M$ is $I$-adically complete if the natural morphism
$M\rightarrow\lim_{\begin{subarray}{c}\longleftarrow\\ n\rightarrow\infty\end{subarray}}M/I^{n}M$ (3.1)
is an isomorphism. In other words, $M$ is complete with respect to $I$-adic topology.
Example 3.2.
We consider the case of $\Lambda_{0}$ and $I$-adic completeness where $I=T\Lambda_{0}$.
  1. 1.
    $\Lambda_{0}$ itself is complete.
  2. 2.
    $\bigoplus_{\mathbb{N}}\Lambda_{0}$ is not complete. The completion is denoted by ^$\widehat{\bigoplus}_{\mathbb{N}}\Lambda_{0}$. Concretely, it consists of a sequence $(x_{i})_{i\in\mathbb{N}}$ of $\Lambda_{0}$ satisfying $\lim_{i\rightarrow\infty}v(x_{i})=\infty$ for the valuation $v$.
  3. 3.
    (^$(\widehat{\bigoplus}_{\mathbb{N}}\Lambda_{0})\otimes_{\Lambda_{0}}(\widehat{\bigoplus}_{\mathbb{N}}\Lambda_{0})$ is not complete. Let us write the basis explicitly: (^$(\widehat{\bigoplus}_{i\in\mathbb{N}}\Lambda_{0}e_{i})\otimes_{\Lambda_{0}}(\widehat{\bigoplus}_{j\in\mathbb{N}}\Lambda_{0}f_{j})$. For example, in the completion, we have $\sum_{i=0}^{\infty}T^{i}e_{i}\otimes f_{i}$, but is not in (^$(\widehat{\bigoplus}_{\mathbb{N}}\Lambda_{0})\otimes_{\Lambda_{0}}(\widehat{\bigoplus}_{\mathbb{N}}\Lambda_{0})$.
    In particular, the category of complete modules is not monoidal with respect to $\otimes_{\Lambda_{0}}$.
  4. 4.
    $\prod_{\mathbb{N}}\Lambda_{0}$ is complete.
  5. 5.
    $\Lambda_{0}[T^{-1}]$ is not complete, since $\Lambda_{0}[T^{-1}]/I^{n}\Lambda_{0}[T^{-1}]=0$ for any $n$.
  6. 6.
    $\Lambda_{0}/\mathfrak{m}$ is complete, since $(\Lambda_{0}/\mathfrak{m})/I^{n}(\Lambda_{0}/\mathfrak{m})=\Lambda_{0}/\mathfrak{m}$.
It is known that the category of complete modules in general does not form an abelian category:
Example 3.3 (Adaptation of [Aut, 110.10]).
We consider the following map
$\varphi\colon\widehat{\bigoplus}_{n\in\mathbb{N}}\Lambda_{0}\rightarrow\prod_{n\in\mathbb{N}}\Lambda_{0},(x_{1},x_{2},x_{3},...)\mapsto(x_{1},Tx_{2},T^{2}x_{3},...).$ (3.2)
in the full subcategory of the complete modules in $\mathrm{Mod}^{\heartsuit}(\Lambda_{0}).$ We would like to check the homomorphism theorem. We first compute the coimage. The coimage is defined by the cokernel of the map $\ker(\varphi)\rightarrow\widehat{\bigoplus}_{n\in\mathbb{N}}\Lambda_{0}$. Hence it is isomorphic to ^$\widehat{\bigoplus}_{n\in\mathbb{N}}\Lambda_{0}$, since $\varphi$ is injective.
On the other hand, the image $\operatorname{im}(\varphi)$ is defined by the kernel of the map $\prod_{n\in\mathbb{N}}\Lambda_{0}\rightarrow\operatorname{coker}(\phi)$. The cokernel is defined by the completion of the cokernel $\operatorname{coker}^{\heartsuit}(\varphi)$ taken in $\mathrm{Mod}^{\heartsuit}(\Lambda_{0})$. Since the element $(1,T,T^{2},..)\in\prod_{\mathbb{N}}\Lambda_{0}$ is not coming from $\varphi$, it defines a nontrivial element in $\operatorname{coker}^{\heartsuit}(\varphi)$. But, for any $n$, the class defined by $(1,T,T^{2},..)$ modulo $\mathfrak{m}^{n}$ is hit by $\varphi$. Hence $(1,T,T^{2},..)$ is zero in the completion. Hence $(1,T,T^{2},...)\in\operatorname{im}(\varphi)$. Hence the canonical morphism $\mathrm{coim}(\varphi)\rightarrow\operatorname{im}(\varphi)$ is not an isomorphism.
As we have seen, the notion of complete modules does not behave well homologically. For this reason, we use the notion of derived complete modules.
Definition 3.4.
Let $A$ be a ring and $I$ is a finitely generated ideal of $A$. We say an object $M\in\mathrm{Mod}(A)$ is derived complete if $\operatorname{Hom}_{\mathrm{Mod}(A)}(A[f^{-1}],M)\cong 0$ for any $f\in I$.
We have the following properties.
Lemma 3.5 ([Ked]).
Let $A$ be a ring and $I$ is a finitely generated ideal of $A$.
  1. 1.
    Any complete module is derived complete.
  2. 2.
    Suppose $M\in\mathrm{Mod}^{\heartsuit}(A)$ is separated i.e., $\bigcap_{n}I^{n}M=0$ and derived complete. Then $M$ is complete.
Definition 3.6.
We denote the subcategory of derived complete modules of $\mathrm{Mod}(\Lambda_{0})$ by $\mathrm{Mod}_{\mathrm{c}}(\Lambda_{0})$. We also set
$\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}}):=\mathfrak{f}^{-1}(\mathrm{Mod}_{\mathrm{c}}(\Lambda_{0})).$ (3.3)
Lemma 3.7.
  1. 1.
    The inclusion $\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}})\subset\mathrm{Mod}(L_{0}^{\mathbb{G}})$ admits a left adjoint. We call it the completion and denote it by $M\mapsto\widehat{M}$. Explicitly, it is given by
    $\widehat{M}:=\lim_{\begin{subarray}{c}\longleftarrow\\ r\rightarrow+\infty\end{subarray}}L_{0}^{\mathbb{G}}/\mathfrak{m}(r)\otimes_{L_{0}^{\mathbb{G}}}M$ (3.4)
  2. 2.
    $\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}})$ is a presentable category. We denote the coproduct (resp. monoidal operation) in $\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}})$ by $\widehat{\bigoplus}$ (resp. $\widehat{\otimes}$).
Proof.
The inclusion is obviously product-preserving, hence we have a left adjoint. The coproduct is given by
^$\widehat{\bigoplus}_{i\in I}\mathcal{E}_{i}:=\widehat{\bigoplus_{i\in I}\mathcal{E}_{i}}.$ (3.5)
Example 3.8.
  1. 1.
    From the above examples, $\Lambda_{0},\prod_{\mathbb{N}}\Lambda_{0},\Lambda_{0}/\mathfrak{m}$ are complete, hence derived complete.
  2. 2.
    The coproduct $\bigoplus_{\mathbb{N}}\Lambda_{0}$ is separated and not complete, hence not derived complete.
  3. 3.
    $\Lambda_{0}[T^{-1}]$ is not complete, not separated. Since $\operatorname{Hom}_{\Lambda_{0}}(\Lambda_{0}[T^{-1}],\Lambda_{0}[T^{-1}])$ is not zero, the module $\Lambda_{0}[T^{-1}]$ is not derived complete.
  4. 4.
    The cokernel of the map
    ^$\widehat{\bigoplus}_{n\in\mathbb{N}}\Lambda_{0}\rightarrow\prod_{n\in\mathbb{N}}\Lambda_{0},(x_{1},x_{2},x_{3},...)\mapsto(x_{1},Tx_{2},T^{2}x_{3},...)$ (3.6)
    in Example 3.3 taken in $\mathrm{Mod}^{\heartsuit}(\Lambda_{0})$ is not complete. But it gives an exact triangle in $\mathrm{Mod}(\Lambda_{0})$, hence derived complete.
4 Almost modules
In the later comparison, we have some discrepancy between sheaf and Novivkov ring which can be ignored by using almost mathematics. In this section, we recall some basic constructions. We refer to [GR03] for general ideas of almost mathematics.
4.1 Almost isomorphism
We have the usual Novikov ring $\Lambda_{0}$ and its maximal ideal $\mathfrak{m}$.
Definition 4.1.
  1. 1.
    For $M\in\mathrm{Mod}(\Lambda_{0})$, we say $M$ is almost zero if $M\otimes_{\Lambda_{0}}\mathfrak{m}=0$.
  2. 2.
    Let $f\colon M\rightarrow N$ be a morphism of $\Lambda_{0}$-modules. We say $f$ is an almost isomorphism if $\ker(f)$ and $\operatorname{coker}(f)$ are almost zero modules.
We note that the full subcategory $\Sigma$ of $\mathrm{Mod}(\Lambda_{0})$ consisting of the almost zero modules is a thick subcategory. We take the quotient $\mathrm{Mod}(\Lambda_{0}^{\mathrm{a}}):=\mathrm{Mod}(\Lambda_{0})/\Sigma$. We denote the quotient functor by
$\mathfrak{a}\colon\mathrm{Mod}(\Lambda_{0})\rightarrow\mathrm{Mod}(\Lambda_{0}^{\mathrm{a}}).$ (4.1)
There exists a right adjoint [GR03] given by
$(-)_{*}:=\operatorname{Hom}_{\mathrm{Mod}(\Lambda_{0}^{\mathrm{a}})}(\Lambda_{0},-)\colon\mathrm{Mod}(\Lambda_{0}^{\mathrm{a}})\rightarrow\mathrm{Mod}(\Lambda_{0});M\mapsto M_{*}.$ (4.2)
We next consider $R:=L_{0}^{\mathbb{G}}$ for some $\mathbb{G}$. We have the forgetful functor $\mathrm{Mod}(L_{0}^{\mathbb{G}})\rightarrow\mathrm{Mod}(\Lambda_{0})$. We set
$\Sigma_{R}:=\mathfrak{f}^{-1}(\Sigma).$ (4.3)
We set
$\mathrm{Mod}(R^{\mathrm{a}}):=\mathrm{Mod}(R)/\Sigma_{R}.$ (4.4)
We denote the quotient functor by
$\mathfrak{a}\colon\mathrm{Mod}(R)\rightarrow\mathrm{Mod}(R^{\mathrm{a}}).$ (4.5)
We say $M,N\in\mathrm{Mod}(R)$ are almost isomorphic if $\mathfrak{a}(M)\cong\mathfrak{a}(N)$.
4.2 Almost isomorphisms
Let $\mathcal{C}$ be a $\Lambda_{0}$-linear category. We denote the $\Lambda_{0}$-linear Yoneda embedding by $\mathcal{Y}\colon\mathcal{C}\rightarrow\mathrm{Mod}(\mathcal{C},\Lambda_{0})$.
Definition 4.2.
  1. 1.
    Let $\mathrm{Mod}(\mathcal{C},\Lambda_{0})$ be the category of $\Lambda_{0}$-modules. An almost zero module $M$ is a module such that $\mathfrak{m}\otimes M\cong 0$. In other words, $M(c)$ is almost zero for any $c\in\mathcal{C}$.
  2. 2.
    We denote the category of almost modules by $\mathrm{Mod}(\mathcal{C},\Lambda_{0}^{\mathrm{a}})$ which is the quotient by the almost zero modules.
  3. 3.
    We say $\mathcal{E},\mathcal{F}\in\mathcal{C}$ are almost isomorphic to $\mathcal{Y}(\mathcal{E})$ and $\mathcal{Y}(\mathcal{F})$ are almost isomorphic. In this case, we denote it by $\mathcal{E}\cong_{\mathrm{a}}\mathcal{F}$.
Lemma 4.3.
Let $f,f^{\prime}\colon\mathcal{E}\rightarrow\mathcal{F}$ be morphisms such that $f-f^{\prime}$ is almost zero. Then $\operatorname{Cone}(f)\cong_{\mathrm{a}}\operatorname{Cone}(f^{\prime})$.
Proof.
By Gabber–Ramero, the module ~$\widetilde{}\mathfrak{m}:=\Lambda_{+}\otimes_{\Lambda_{0}}\Lambda_{+}$ is flat. We then have an almost isomorphism
~$\widetilde{}\mathfrak{m}\otimes_{\Lambda_{0}}\mathcal{Y}(\operatorname{Cone}(f^{\prime}))\rightarrow\mathcal{Y}(\operatorname{Cone}(f)).$ (4.6)
We have
$\mathcal{Y}(\operatorname{Cone}(f^{\prime}))\cong_{\mathrm{a}}\widetilde{}\mathfrak{m}\otimes_{\Lambda_{0}}\mathcal{Y}(\operatorname{Cone}(f^{\prime}))\cong_{\mathrm{a}}\mathcal{Y}(\operatorname{Cone}(f)).$ (4.7)
For general $\mathbb{G}$, we slightly modify the setup:
Definition 4.4.
Let $\mathbb{G}$ be a subgroup of $\mathbb{R}$. A category over $L_{0}^{\mathbb{G}}$ is a tuple
  1. 1.
    A category $\mathcal{C}$, and
  2. 2.
    a group homomorphism $T_{\bullet}\colon\mathbb{R}/\mathbb{G}\rightarrow\operatorname{Aut}(\mathcal{C})$.
with a homomorphism
$L_{0}^{\mathbb{G}}\rightarrow\operatorname{End}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}).$ (4.8)
Let $\mathcal{C}$ be a category over $L_{0}^{\mathbb{G}}$. Then we have a functor
$\mathcal{Y}\colon\mathcal{C}\rightarrow\mathrm{Mod}(\mathcal{C},L_{0}^{\mathbb{G}}):=\mathop{\mathrm{\bf{Fun}}}(\mathcal{C}^{op},\mathrm{Mod}(L_{0}^{\mathbb{G}}));\mathcal{E}\mapsto\operatorname{Hom}(-,\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}\mathcal{E}).$ (4.9)
Definition 4.5.
  1. 1.
    An almost zero module $M\in\mathrm{Mod}(\mathcal{C},L_{0}^{\mathbb{G}})$ is a module such that $M(c)\in\Sigma_{L_{0}^{\mathbb{G}}}$ for any $c\in\mathcal{C}$.
  2. 2.
    We denote the category of almost modules by $\mathrm{Mod}(\mathcal{C},{L_{0}^{\mathbb{G}}}^{\mathrm{a}})$ which is the quotient by the almost zero modules.
  3. 3.
    We say $\mathcal{E},\mathcal{F}\in\mathcal{C}$ are almost isomorphic to $\mathcal{Y}(\mathcal{E})$ and $\mathcal{Y}(\mathcal{F})$ are almost isomorphic. In this case, we denote it by $\mathcal{E}\cong_{\mathrm{a}}\mathcal{F}$.
Similarly, we can prove that the extensions by almost same morphisms are almost isomorphic.
4.3 Almost equivalence
Definition 4.6.
Let $\mathcal{C}_{1},\mathcal{C}_{2}$ be categories defined over $\Lambda_{0}$. Let $F\colon\mathcal{C}_{1}\rightarrow\mathcal{C}_{2}$ be a $\Lambda_{0}$-linear functor. We say $F$ is an almost equivalence if it satisfies the following:
  1. 1.
    For any $c,c^{\prime}\in\mathcal{C}_{1}$, the induced morphism
    $\operatorname{Hom}_{\mathcal{C}_{1}}(c,c^{\prime})\rightarrow\operatorname{Hom}_{\mathcal{C}_{2}}(F(c),F(c^{\prime}))$ (4.10)
    is an almost isomorphism.
  2. 2.
    For any $c^{\prime}\in\mathcal{C}_{2}$, there exists $c\in\mathcal{C}_{1}$ such that $F(c)$ is almost isomorphic to $c^{\prime}$.
In the following, we give a little generalization of the above notion:
Definition 4.7.
Let $\mathcal{C}_{1},\mathcal{C}_{2}$ be categories over $L_{0}^{\mathbb{G}}$.
  1. 1.
    A morphism $F\colon\mathcal{C}_{1}\rightarrow\mathcal{C}_{2}$ is a functor from $\mathcal{C}_{1}$ to $\mathcal{C}_{2}$ commuting with $T_{\bullet}$.
  2. 2.
    We say a morphism $F\colon\mathcal{C}_{1}\rightarrow\mathcal{C}_{2}$ is almost fully faithful (, or almost embedding, $\mathcal{C}_{1}\hookrightarrow_{\mathrm{a}}\mathcal{C}_{2}$) if the following holds: For any $\alpha,\alpha^{\prime}\in\mathcal{C}_{1}$, the induced morphism
    $\operatorname{Hom}_{\mathcal{C}_{1}}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}\alpha,\alpha^{\prime})\rightarrow\operatorname{Hom}_{\mathcal{C}_{2}}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}F(\alpha),F(\alpha^{\prime}))$ (4.11)
    is an almost isomorphism in $\mathrm{Mod}(L_{0}^{\mathbb{G}})$.
  3. 3.
    We say a morphism $F\colon\mathcal{C}_{1}\rightarrow\mathcal{C}_{2}$ is almost essentially surjective if the following holds: For any $\alpha^{\prime}\in\mathcal{C}_{2}$, there exists $\alpha\in\mathcal{C}_{1}$ such that $F(\alpha)$ is almost isomorphic to $c^{\prime}$.
  4. 4.
    We say $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are almost equivalent (, or $\mathcal{C}_{1}\cong_{\mathrm{a}}\mathcal{C}_{2}$) if there exists a morphism from $\mathcal{C}_{1}$ to $\mathcal{C}_{2}$ such that $f$ is almost fully faithful and almost essentially surjective.
We will deal with the following two examples:
Example 4.8.
We consider the category $\mathrm{Mod}(L_{0}^{\mathbb{G}})$. Since an object of $\mathrm{Mod}(L_{0}^{\mathbb{G}})$ carries a $\mathbb{R}/\mathbb{G}$-grading, we can shift it. We denote the resulting shift functor by $T_{c}$. Then, for any $M,N\in\mathrm{Mod}(L_{0}^{\mathbb{G}})$, we have
$\operatorname{Hom}_{\mathrm{Mod}(L_{0}^{\mathbb{G}})}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}M,N),$ (4.12)
which is an $L_{0}^{\mathbb{G}}$-module.
Example 4.9.
The category $\mu^{\mathbb{G}}(M)$ will be introduced in the next section. We have shift operations $T_{c}$ parmetrized by $c\in\mathbb{R}/\mathbb{G}$. Then, for any $\mathcal{E},\mathcal{F}\in\mu^{\mathbb{G}}(M)$, we have
$\operatorname{Hom}_{\mu^{\mathbb{G}}(M)}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}\mathcal{E},\mathcal{F}),$ (4.13)
which is an $L_{0}^{\mathbb{G}}$-module.
5 Equivariant sheaves and the Novikov ring
In this section and the next section, we relate Tamarkin categories with Novikov rings precisely.
5.1 Basics
Let $\mathbb{R}_{t}$ be the $1$-dimensional real vector space with the standard coordinate $t$.
We consider the addition action of a subgroup $\mathbb{G}\subset\mathbb{R}$ on $\mathbb{R}_{t}$ as a discrete group action. Then we can consider the derived category of equivariant $\mathbb{K}$-module sheaves $\mathrm{Sh}^{\mathbb{G}}(\mathbb{R}_{t},\mathbb{K})$.
We denote the subcategory spanned by the object whose microsupport contained in ×$\mathbb{R}\times\mathbb{R}_{\leq 0}\subset\mathbb{R}\times\mathbb{R}\cong T^{*}\mathbb{R}_{t}$ by $\mathrm{Sh}^{\mathbb{G}}_{\mathbb{R}_{\leq 0}}(\mathbb{R}_{t},\mathbb{K})$. We set
$\mu^{\mathbb{G}}(*):=\mathrm{Sh}^{\mathbb{G}}_{\mathbb{R}_{>0}}(\mathbb{R}_{t},\mathbb{K}):=\mathrm{Sh}^{\mathbb{G}}(\mathbb{R}_{t},\mathbb{K})/\mathrm{Sh}^{\mathbb{G}}_{\mathbb{R}_{\leq 0}}(\mathbb{R}_{t},\mathbb{K}).$ (5.1)
We have the following:
Lemma 5.1 ([Kuwal, Kuw22]).
  1. 1.
    $\mu^{\mathbb{G}}(*)$ has a monoidal structure defined by the convolution product.
  2. 2.
    We equip the sheaf $1_{\mu}:=\bigoplus_{c\in\mathbb{G}}\mathbb{K}_{t\geq c}$ with an obvious equivariant structure. Then it defines an object of $\mu^{\mathbb{G}}(*)$ and is a monoidal unit.
  3. 3.
    We have $H^{0}\operatorname{End}_{\mu^{\mathbb{G}}(*)}(1_{\mu})\cong\Lambda_{0}^{\mathbb{G}}$. As a corollary of 2 and 3, $\mu^{\mathbb{G}}(*)$ is enriched over $\Lambda_{0}^{\mathbb{G}}$.
We strengthen the result a little more.
Lemma 5.2.
We have an almost isomorphism of almost $L_{0}^{\mathbb{G}}$-modules
$\operatorname{End}_{\mu^{\mathbb{G}}(*)}(1_{\mu})\cong_{\mathrm{a}}\Lambda_{0}^{\mathbb{G}}.$ (5.2)
More precisely, the higher cohomologies of $\operatorname{End}_{\mu^{\mathbb{G}}(*)}(1_{\mu})$ are almost zero, but not zero.
Proof.
For the case when $\mathbb{G}=\{0\}$ and $\mathbb{G}\cong\mathbb{Z}$, there are no higher cohomologies. In the following, we only prove the case when $\mathbb{G}=\mathbb{R}$. Other cases (i.e., dense subgroups of $\mathbb{R}$) can be proved similarly.
We first consider the following exact triangle:
$\operatorname{Hom}(\mathbb{K}_{t\geq 0},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\rightarrow\operatorname{Hom}(\mathbb{K}_{\mathbb{R}},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\rightarrow\operatorname{Hom}(\mathbb{K}_{t<0},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\rightarrow.$ (5.3)
Since the $i$-th cohomologies vanish for $i>1$, we will only take care of $H^{1}\operatorname{Hom}(\mathbb{K}_{\mathbb{R}},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})$. (Note that $H^{1}\operatorname{Hom}(\mathbb{K}_{\mathbb{R}},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\cong H^{1}\operatorname{Hom}(\mathbb{K}_{t<0},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})$.)
Now we consider the exact triangle:
$\operatorname{Hom}(\mathbb{K}_{\mathbb{R}},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t<c})\rightarrow\operatorname{Hom}(\mathbb{K}_{\mathbb{R}},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{\mathbb{R}})\rightarrow\operatorname{Hom}(\mathbb{K}_{\mathbb{R}},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\rightarrow$ (5.4)
Note that
$\operatorname{Hom}(\mathbb{K}_{\mathbb{R}},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{\mathbb{R}})\simeq\bigoplus_{c\in\mathbb{R}}\mathbb{K}.$ (5.5)
Hence we have $H^{1}\operatorname{Hom}(\mathbb{K}_{\mathbb{R}},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\simeq 0$. Hence
$H^{1}\operatorname{Hom}(\mathbb{K}_{t\geq 0},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\simeq\operatorname{coker}(H^{0}\operatorname{Hom}(\mathbb{K}_{\mathbb{R}},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\rightarrow H^{0}\operatorname{Hom}(\mathbb{K}_{t<0},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})).$ (5.6)
Note that the right hand side is isomorphic to
$\operatorname{coker}(H^{0}\operatorname{Hom}(\mathbb{K}_{-a<t},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\rightarrow H^{0}\operatorname{Hom}(\mathbb{K}_{-a<t<0},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c}))$ (5.7)
for any $a>0$. This implies that the action of $T^{a}$ on (5.6) is zero for any $a>0$. Hence it is almost zero. This completes the proof. ∎
We further have the following:
Lemma 5.3.
We have an almost isomorphism of almost $L_{0}^{\mathbb{G}}$-modules
$\operatorname{End}_{\mu^{\mathbb{G}}(*)}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu})\cong L_{0}^{\mathbb{G}}.$ (5.8)
Proof.
By using the preceding lemma, we have a sequence of almost isomorphisms
$\begin{split}\operatorname{End}_{\mu^{\mathbb{G}}(*)}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu})\cong\prod_{c\in\mathbb{R}/\mathbb{G}}\operatorname{Hom}_{\mu^{\mathbb{O}}(*)}(\mathbb{K}_{t\geq 0},\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\cong\prod_{c\in\mathbb{R}/\mathbb{G}}\Lambda_{0}\cong L_{0}^{\mathbb{G}}\end{split}$ (5.9)
By Lemma 2.6, we get an isomorphism of the underlying $\mathbb{K}$-modules. One can see that this isomorphism preserves the algebra structure. ∎
5.2 Derived completeness
Lemma 5.4.
For any $\mathcal{E},\mathcal{F}\in\mu^{\mathbb{G}}(*)$, we have $\operatorname{Hom}_{\mu^{\mathbb{G}}(*)}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}\mathcal{E},\mathcal{F})\in\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}})$.
Proof.
We will show the space $\operatorname{Hom}_{\mu^{\mathbb{R}}(*)}(\mathcal{E},\mathcal{F})$ for $\mathcal{E},\mathcal{F}$ is derived complete. The other cases will follow from similar arguments. In other words, it is enough to show the homotopy limit of the sequence
[…]
is zero [Lur, Cor 4.2.8]. For the notation, we denote it by
$\lim_{\begin{subarray}{c}\longleftarrow\\ i\rightarrow\infty\end{subarray}}\operatorname{Hom}(\mathcal{E},\mathcal{F}).$ (5.11)
By using the internal hom defined in [Tam18, Kuwal], we have
$\begin{split}\lim_{\begin{subarray}{c}\longleftarrow\\ i\rightarrow\infty\end{subarray}}\operatorname{Hom}_{\mu^{\mathbb{R}}(*)}(\mathcal{E},\mathcal{F})&=\operatorname{Hom}_{\mu^{\mathbb{R}}(*)}(\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c},\lim_{\begin{subarray}{c}\longleftarrow\\ i\rightarrow\infty\end{subarray}}\mathop{{\mathcal{H}}om}\nolimits^{\star_{\mathbb{R}}}(\mathcal{E},\mathcal{F}))\\ &=\operatorname{Hom}_{\mu^{\mathbb{O}}(*)}(\mathbb{K}_{t\geq 0},\lim_{\begin{subarray}{c}\longleftarrow\\ i\rightarrow\infty\end{subarray}}\mathop{{\mathcal{H}}om}\nolimits^{\star_{\mathbb{R}}}(\mathcal{E},\mathcal{F}))\\ \end{split}$ (5.12)
We set $\mathop{{\mathcal{H}}om}\nolimits^{\star_{\mathbb{G}}}(\mathcal{E},\mathcal{F})=:\mathcal{G}$. Here ${\lim_{\begin{subarray}{c}\longleftarrow\\ i\rightarrow\infty\end{subarray}}\mathcal{G}}$ is the homotopy limit of the sequence
[…]
We then have
$\begin{split}\Gamma(\mathbb{K}_{\left\{t\geq 0\right\}},\lim_{\begin{subarray}{c}\longleftarrow\\ i\rightarrow\infty\end{subarray}}\mathcal{G})&\cong\Gamma(\mathbb{K}_{\left\{t\geq 0\right\}},\lim_{\begin{subarray}{c}\longleftarrow\\ i\rightarrow\infty\end{subarray}}\mathcal{G})\\ &\cong\lim_{\begin{subarray}{c}\longleftarrow\\ i\rightarrow\infty\end{subarray}}\Gamma(\mathbb{K}_{\left\{t\geq 0\right\}},\mathcal{G})\\ &\cong\lim_{\begin{subarray}{c}\longleftarrow\\ c\rightarrow\infty\end{subarray}}\Gamma(\mathbb{K}_{\left\{t\geq c\right\}},\mathcal{G})\\ &\cong\Gamma(\lim_{\begin{subarray}{c}\longrightarrow\\ c\rightarrow\infty\end{subarray}}\mathbb{K}_{\left\{t\geq c\right\}},\mathcal{G})\cong 0.\end{split}$ (5.14)
This completes the proof. ∎
5.3 Morita functor
By Lemma 5.4, we have the Morita functor
$\mathfrak{A}\colon\mu^{\mathbb{G}}(*)\rightarrow\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}});\mathcal{E}\mapsto\operatorname{Hom}_{\mu^{\mathbb{G}}(*)}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu},\mathcal{E}).$ (5.15)
Theorem 5.5.
The functor $\mathfrak{A}$ is an almost equivalence.
Proof of Theorem 5.5.
Step 1
We first show the conservativity of the functor: If an object $\mu^{\mathbb{G}}(T^{*}\{*\})$ satisfies $\operatorname{Hom}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu},\mathcal{E})=0$, then $\mathcal{E}=0$. Actually, if $\mathcal{E}$ satisfies $\operatorname{Hom}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu},\mathcal{E})=0$, it means the vanishing of positive microsupport, which means $\mathcal{E}=0$.
Step 2
We next prove the following:
Claim 5.6.
$\operatorname{Hom}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu},\bigoplus_{i\in I}T_{d_{i}}1_{\mu})\cong_{\mathrm{a}}\widehat{\bigoplus}_{i\in I}\operatorname{Hom}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu},T_{d_{i}}1_{\mu})$ (5.16)
in $\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}})$.
Proof of claim.
We consider the case of $\mathbb{G}=\mathbb{R}$. Other cases are similar.
We first prove
$\operatorname{Hom}(T_{c}1_{\mu},\bigoplus_{i\in I}T_{d_{i}}1_{\mu})\cong\widehat{\bigoplus}_{i\in I}\operatorname{Hom}(T_{c}1_{\mu},T_{d_{i}}1_{\mu}).$ (5.17)
We only consider the case when $c=0$, since other cases are similar. We first replace the left hand side with
$\operatorname{Hom}(1_{\mu},\bigoplus_{i\in I}1_{\mu})\cong\operatorname{Hom}(\mathbb{K}_{t\geq 0},\bigoplus_{I}\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c}).$ (5.18)
Then we have
$\operatorname{Hom}(\mathbb{K}_{t\geq 0},\bigoplus_{I}\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\rightarrow\operatorname{Hom}(\mathbb{K},\bigoplus_{I}\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\rightarrow\operatorname{Hom}(\mathbb{K}_{(-\infty,0)},\bigoplus_{I}\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\rightarrow.$ (5.19)
As in the proof of Lemma 5.2, we can see that $H^{1}(\operatorname{Hom}(\mathbb{K},\bigoplus_{I}\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c}))\simeq 0$. Then one can go to the cohomology exact sequence as
$0\rightarrow H^{0}\operatorname{Hom}(\mathbb{K}_{t\geq 0},\bigoplus_{I}\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\rightarrow\widehat{\bigoplus}_{c\in(-\infty,+\infty)}\mathbb{K}\rightarrow\widehat{\bigoplus}_{c\in(-\infty,0)}\mathbb{K}\rightarrow H^{1}\operatorname{Hom}(\mathbb{K}_{t\geq 0},\bigoplus_{I}\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\rightarrow 0$ (5.20)
Here the completion of ^$\widehat{\bigoplus}_{c\in(-\infty,0)}\mathbb{K}$ is taken in the direction to $\infty$- ∞ and 00 and the completion of ^$\widehat{\bigoplus}_{c\in(-\infty,0)}\mathbb{K}$ is taken in the direction to $\infty$- ∞ and +$\infty$+ ∞. Hence we conclude that
$H^{i}\operatorname{Hom}(\mathbb{K}_{t\geq 0},\bigoplus_{I}\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\cong\begin{cases}&\Lambda_{0}\text{ if $i=0$}\\ &\operatorname{coker}\left(\bigoplus_{[1,0)}\mathbb{K}\rightarrow\widehat{\bigoplus}_{[1,0)}\mathbb{K}\right)\text{ if $i=1$}\\ &0\text{ otherwise}.\end{cases}$ (5.21)
where the morphism in the second line is a natural one. Then the almostization kill the degree one morphisms.
By (5.17), we have
$\operatorname{Hom}(1_{\mu},\bigoplus_{i\in I}T_{d_{i}}1_{\mu})\cong\widehat{\bigoplus}_{i\in I}\operatorname{Hom}(T_{c}1_{\mu},T_{d_{i}}1_{\mu}).$ (5.22)
This completes the proof. ∎
Step 3
By using the standard argument (cf. [SS03]) and Step 2, for any object $\mathcal{E}\in\mu^{\mathbb{G}}(*)$, we can construct an exact triangle
$A\rightarrow\mathcal{E}\rightarrow B\xrightarrow{[1]}$ (5.23)
where $A$ is in the colimits of $\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu}$ and $B$ is in the right orthogonal of such colimits. In Step 1, we check that such an orthogonal is zero. Hence any $\mathcal{E}$ is in the colimits of $\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu}$. Since $\mathfrak{A}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu})\cong_{\mathrm{a}}L_{0}^{\mathbb{G}}$ by Lemma 5.3 and
$\operatorname{End}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu})\cong_{\mathrm{a}}\operatorname{End}(L_{0}^{\mathbb{G}})=L_{0}^{\mathbb{G}},$ (5.24)
we conclude that the functor $\mathfrak{A}$ is almost fully faithful.
Also, since $L_{0}^{\mathbb{G}}$ is a generator of $\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}})$ and $\mathfrak{A}$ is cocontinuous, the almost essential surjectivity follows. This completes the proof.
6 Global version
6.1 Reminders on the Lurie tensor product
We follow Volpe’s exposition [Vol21]. Let $Cat_{comp}$ be the category of the cocomplete categories and the morphisms are cocontinuous functors. For cocomplete categories $\mathcal{C},\mathcal{D}$, there exists a cocomplete category $\mathcal{C}\otimes_{L}\mathcal{D}$ with a functor $\mathcal{C}\times\mathcal{D}\rightarrow\mathcal{C}\otimes_{L}\mathcal{D}$ satisfying
$Fun(\mathcal{C}\times\mathcal{D},\mathcal{E})\cong Fun(\mathcal{C}\otimes_{L}\mathcal{D},\mathcal{E})$ (6.1)
where the left hand side denotes the functors preserving variable-wise colimits and the right hand side denotes the functors preserving colimits. The resulting category $\mathcal{C}\otimes_{L}\mathcal{D}$ is called Lurie’s tensor product. In the following, we simply denote $\otimes_{L}$ by $\otimes$.
We will mainly use the following properties.
Lemma 6.1 ([Vol21, Corollary 2.30]).
Let $M$ and $N$ be manifolds.
  1. 1.
    We have an equivalence. $\mathrm{Sh}(M,\mathbb{K})\otimes\mathrm{Sh}(N,\mathbb{K})\cong\mathrm{Sh}(M\times N,\mathbb{K})$.
  2. 2.
    Let $\mathcal{C}$ be a presentable $\mathbb{K}$-linear category. Then we have $\mathrm{Sh}(M,\mathbb{K})\otimes\mathcal{C}\cong\mathrm{Sh}(M,\mathcal{C})$.
6.2 Tamarkin-type category
Let $M$ be a manifold. We consider the category $\mathrm{Sh}^{\mathbb{G}}(M\times\mathbb{R}_{t},\mathbb{K})$ of the equivariant $\mathbb{K}$-module sheaves on $M\times\mathbb{R}_{t}$ with respect to the discrete $\mathbb{G}$-action on the left component. We denote the subcategory spanned by the object whose microsupport contained in $T^{*}M\times\mathbb{R}\times\mathbb{R}_{\leq 0}$ by $\mathrm{Sh}^{\mathbb{G}}_{\leq 0}(M\times\mathbb{R}_{t},\mathbb{K})$. We set
$\mu^{\mathbb{G}}(T^{*}M):=\mathrm{Sh}^{\mathbb{G}}_{>0}(M\times\mathbb{R}_{t},\mathbb{K}):=\mathrm{Sh}^{\mathbb{G}}(M\times\mathbb{R}_{t},\mathbb{K})/\mathrm{Sh}^{\mathbb{G}}_{\leq 0}(M\times\mathbb{R}_{t},\mathbb{K}).$ (6.2)
Then $\mu^{\mathbb{G}}(T^{*}M)$ is defined over $\mathrm{Mod}(L_{0}^{\mathbb{G}})$.
The following is known (cf. [IK, KSZ]):
Lemma 6.2 (Lurie tensor product).
$\mu^{\mathbb{G}}(T^{*}M)\cong\mathrm{Sh}(M,\mathbb{K})\otimes\mu^{\mathbb{G}}(*).$ (6.3)
By using Theorem 5.5 and Lemma 6.1, we further have
$\mu^{\mathbb{G}}(T^{*}M)\cong_{\mathrm{a}}\mathrm{Sh}(M,\mathbb{K})\otimes\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}})\cong\mathrm{Sh}(M,\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}})).$ (6.4)
Since $\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}})\hookrightarrow\mathrm{Mod}(L_{0}^{\mathbb{G}})$ is a right adjoint, it induces
$\mathrm{Sh}(M,\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}}))\hookrightarrow\mathrm{Sh}(M,\mathrm{Mod}(L_{0}^{\mathbb{G}})).$ (6.5)
Since $M$ is a manifold, the category $\mathrm{Sh}(M,\mathrm{Mod}(L_{0}^{\mathbb{G}}))$ gets identified with the derived category $\mathrm{Sh}(M,L_{0}^{\mathbb{G}})$ of sheaves on $M$ whose values are $L_{0}^{\mathbb{G}}$-modules.
Corollary 6.3.
We have an almost embedding:
$\mu^{\mathbb{G}}(T^{*}M)\hookrightarrow_{\mathrm{a}}\mathrm{Sh}(M,L_{0}^{\mathbb{G}}).$ (6.6)
6.3 Sheaf quantizations
Let $M$ be a manifold. We denote the cotangent coordinate of $\mathbb{R}_{t}$ by $\tau$. We set
$\left\{\tau>0\right\}:=\left\{(p,(t,\tau))\in T^{*}M\times T^{*}\mathbb{R}_{t}\mathrel{}\middle|\mathrel{}\tau>0\right\}.$ (6.7)
[…]
Definition 6.4 (Sheaf quantization).
An object of $\mu^{\mathbb{G}}(T^{*}M)$ is a sheaf quantization of a Lagrangian submanifold $L$ if
  1. 1.
    $\mu\mathrm{supp}(\mathcal{E})=L$, and
  2. 2.
    the microstalks are finite dimensional.
For the construction and properties of sheaf quantizations, see the companion paper [IK]. We denote the category of $\mu^{\mathbb{G}}(T^{*}M)$ consisting of sheaf quantizations of projection-finite end-conic Lagrangians by $\mathrm{SQ}^{\mathbb{G}}(T^{*}M)$.
Corollary 6.5.
We have an almost embedding
$\mathrm{SQ}^{\mathbb{G}}(T^{*}M)\hookrightarrow_{\mathrm{a}}\mathrm{Sh}(M,L_{0}^{\mathbb{G}}).$ (6.9)
6.4 Variant 1: Liouville manifold
We can easily generalize the construction to the case of Liouville manifolds. Let $X$ be a Liouville manifold. We denote the category of microsheaves over $X$ by $\mu sh(X)$. We set $\mu^{\mathbb{G}}(X):=\mu sh(X)\otimes\mu^{\mathbb{G}}(*)$. We then have $\mu^{\mathbb{G}}(X)\cong_{\mathrm{a}}\mu sh(X)\otimes\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}})$.
6.5 Variant 2: Energy cutoff
We sometime would like to discuss the energy cutoff setup.
Let $N$ be a manifold and $\mathbb{R}_{s<a}:=(-\infty,a)$ for $a>0$. Then we run the above theory to get $\mu^{\mathbb{G}}(T^{*}N\times T^{*}\mathbb{R}_{s<a})$. We consider the subcategory spanned by doubling movies $\mu^{\mathbb{G}}_{<a}(T^{*}N)$: We say an object $\mathcal{E}\in\mathrm{Sh}^{\mathbb{G}}(N\times\mathbb{R}_{s}\times\mathbb{R}_{t})$ is a doubling movie if it satisfies
[…]
for some $A\subset T^{*}M\times\mathbb{R}_{t}\times\mathbb{R}_{\tau>0}$ where
$\begin{split}AA:=&\left\{(p,s,0)\in(T^{*}M\times T^{*}\mathbb{R}_{t})\times T^{*}\mathbb{R}_{s<c}\mathrel{}\middle|\mathrel{}p\in A,s\geq 0\right\}\\ &\cup\left\{(p^{\prime},t,\tau,s,\sigma)\in T^{*}M\times T^{*}\mathbb{R}_{t}\times T^{*}\mathbb{R}_{s<c}\mathrel{}\middle|\mathrel{}(p^{\prime},t-s,\tau)\in A,s\geq 0,\tau=-\sigma\right\}.\end{split}$ (6.11)
We then have
$\mu^{\mathbb{G}}_{<a}(T^{*}N)\cong\mathrm{Sh}(N,\mathbb{K})\otimes\mu^{\mathbb{G}}_{<a}(*).$ (6.12)
Lemma 6.6.
$\mathrm{Mod}(L_{0}^{\mathbb{G}}/T^{a}L_{0}^{\mathbb{G}})\hookrightarrow_{\mathrm{a}}\mu^{\mathbb{G}}_{<a}(*)$ (6.13)
Proof.
We set
$\mu^{\mathbb{G}}_{<a}(*)\ni 1_{\mu,a}:=\bigoplus_{c\in\mathbb{G}}\mathbb{K}_{\left\{(s,t)\mathrel{}\middle|\mathrel{}0<s<a,c\leq t<s+c\right\}}.$ (6.14)
The endmorphism of this can be easily computed as almost $L_{0}^{\mathbb{G}}/T^{a}L_{0}^{\mathbb{G}}$. We then have a functor
$\mathrm{Mod}(L_{0}^{\mathbb{G}}/T^{a}L_{0}^{\mathbb{G}})\rightarrow\mu^{\mathbb{G}}_{<a}(*);M\mapsto M\otimes 1_{\mu,a}.$ (6.15)
Then this is almost fully faithful. ∎
Remark 6.7.
The functor is not essentially surjective. We give an example which are not in the essential image. Consider the following nonzero infinite complex
$\cdots\xrightarrow{T^{1/2}}1_{\mu,1}\xrightarrow{T^{1/2}}1_{\mu,1}\xrightarrow{T^{1/2}}1_{\mu,1}\xrightarrow{T^{1/2}}\cdots$ (6.16)
One can see that this is orthogonal to $1_{\mu,1}$. hence $1_{\mu,1}$ does not generate the whole category. The author would like to thank T. Asano for asking a related question on the draft version.
We immediately have the following.
Corollary 6.8.
$\mu^{\mathbb{G}}_{<a}(T^{*}N)\hookleftarrow_{\mathrm{a}}\mathrm{Sh}(N,\mathrm{Mod}(L_{0}^{\mathbb{G}}/T^{a}L_{0}^{\mathbb{G}})).$ (6.17)
In particular, if $\mathbb{G}=\mathbb{R}$, we have
$\mu^{\mathbb{R}}_{<a}(T^{*}N)\hookleftarrow_{\mathrm{a}}\mathrm{Sh}(N,\Lambda_{0}/T^{a}\Lambda_{0}).$ (6.18)
6.6 Variant 3: Higher-dimensional version
For the use of $\hbar$-Riemann–Hilbert correspondence, we would like to mention the following version:
Let $\gamma$ be a simplicial closed polyhedral proper convex cone in $\mathbb{R}^{n}$ with nonempty interior. Then $\gamma$ has a semigroup structure with respect to the addition. We denote the corresponding polynomial ring by $\mathbb{K}[\gamma]$. We denote the indeterminate corresponding to $a\in\gamma$ by $T^{a}$. Let ||$\cdot$|| ⋅ | be the Euclidean norm of $\mathbb{R}^{n}$. For $r\in\mathbb{R}_{>0}$, we denote the ideal of $\mathbb{K}[\gamma]$ generated by $T^{a}$’s with $|a|>r$ by $\mathfrak{m}(r)$. We set
$\Lambda_{0}^{\gamma}:=\lim_{\begin{subarray}{c}\longleftarrow\\ r\rightarrow\infty\end{subarray}}\mathbb{K}[\gamma\cap\mathbb{G}]/\mathfrak{m}(r).$ (6.19)
We say that $\gamma$ is simplicial if there exists a linear isomorphism of $\mathbb{R}^{n}$ which gives an isomorphism $\gamma\cong\mathbb{R}^{n}_{\geq 0}$.
We consider $\mathbb{R}^{n}$-equivariant sheaves on $M\times\mathbb{R}^{n}$. We denote the subcategory spanned by the object whose microsupport contained in $T^{*}M\times\mathbb{R}^{n}\times(-\gamma^{\vee})$ by $\mathrm{Sh}^{\mathbb{R}^{n}}_{\leq 0}(M\times\mathbb{R}_{t},\mathbb{K})$ where $\vee$ is polar dual. We set
$\mu^{\gamma}(T^{*}M):=\mathrm{Sh}^{\mathbb{R}^{n}}_{\mathrm{Int}(\gamma^{\vee})}(M\times\mathbb{R}_{t},\mathbb{K}):=\mathrm{Sh}^{\mathbb{R}^{n}}(M\times\mathbb{R}_{t},\mathbb{K})/\mathrm{Sh}^{\mathbb{R}^{n}}_{(-\gamma^{\vee})}(M\times\mathbb{R}_{t},\mathbb{K}).$ (6.20)
By tensoring Theorem 6.3 $n$ times, we ge the following:
Theorem 6.9.
Suppose $\gamma$ is simplicial. We have an almost embedding
$\mu^{\gamma}(T^{*}M)\hookrightarrow_{\mathrm{a}}\mathrm{Sh}(M,\Lambda_{0}^{\gamma}).$ (6.21)
Remark 6.10.
If $n=2$, every $\gamma$ is simplicial. It should be possible to remove the restriction on $\gamma$ and $\mathbb{G}$. But, so far, we don’t know any application of such a generalization.
7 Application I: Non-conic microlocal sheaf theory
By Theorem 6.3, one can imagine that microlocal sheaf theory over Novikov rings are non-conic. In this section, we develop such a theory.
7.1 Non-conic microsupport
Let $R$ be a real valuation ring. Our examples are $\Lambda_{0}^{\mathbb{G}}$ for a dense subgroup $\mathbb{G}\subset\mathbb{R}$. We denote the valuation by $v$. We set
$R_{c}:=R/\left\{r\in R\mathrel{}\middle|\mathrel{}v(r)>c\right\}.$ (7.1)
Let $U$ be an open subset of $M$. Let $\phi$ be a continuous function on $U$ which is bounded below. For any connected open subset $V\subset U$, we denote the infimum value of $\phi$ by $\phi_{V}$. We define a sheaf $R^{\phi}$ as follows: For any connected open subset $V$, we set $R^{\phi}(V):=R$. For an open inclusion of connected open subsets $W\subset V$, we have $\phi_{W}\geq\phi_{V}$, we set a structure morphism $R^{\phi}(V)\rightarrow R^{\phi}(W)$ by $T^{\phi_{W}-\phi_{V}}$. Sheafifying this, we get a sheaf on $U$. We also set $R^{\phi}_{c}:=R^{\phi}\otimes_{R}R_{c}$ for $c\geq 0$.
Definition 7.1.
For $\mathcal{E}\in\mathrm{Mod}(R_{M})$, we set $\mu\mathrm{supp}(\mathcal{E})$ to be the closure of the complement of the following set:
$\left\{(x,\xi)\mathrel{}\middle|\mathrel{}\mathop{{\mathcal{H}}om}\nolimits(R^{\phi}_{c},\mathcal{E})_{x}\simeq 0\text{ for any $c\geq 0$ and $C^{1}$-function $\phi$ with $d\phi(x)=\xi$}\right\}.$ (7.2)
The followings are obvious from the definition.
Lemma 7.2.
  1. 1.
    $\mu\mathrm{supp}(\mathcal{E})$ is closed.
  2. 2.
    Let $\mathcal{E}_{1}\rightarrow\mathcal{E}_{2}\rightarrow\mathcal{E}_{3}\xrightarrow{[1]}$ be an exact triangle. Then we have
    $\mu\mathrm{supp}(\mathcal{E}_{2})\subset\mu\mathrm{supp}(\mathcal{E}_{1})\cup\mu\mathrm{supp}(\mathcal{E}_{3}).$ (7.3)
Although we do not address here, generalizing functoriality results of microsupport in [KS94] to our setup should be an interesting problem.
7.2 Relation to usual microsupport
Since an object of $\mathrm{Mod}(R_{M})$ is a sheaf on a manifold, we can also define the usual microsupport of [KS94].
Definition 7.3.
For a subset $A\subset T^{*}M$, we set
$\mathbb{R}_{\geq 0}\cdot A:=\left\{(x,\xi)\in T^{*}M\mathrel{}\middle|\mathrel{}(x,\xi^{\prime})\in A,c\in\mathbb{R}_{\geq 0}\text{ s.t.}(x,\xi)=(x,c\xi^{\prime})\right\}$ (7.4)
Note that the usual microsupport does not have much information for our sheaves, as the following proposition suggests.
Proposition 7.4.
For $\mathcal{E}\in\mathrm{Mod}(R_{M})$, we have >$SS\mathbb{R}_{>0}\cdot\mu\mathrm{supp}(\mathcal{E})\subset\SS(\mathcal{E})$.
7.3 Relation to non-conic microsupport
Proposition 7.5.
For an object $\mathcal{E}\in\mu^{\mathbb{G}}(T^{*}M)$, we have
$\mu\mathrm{supp}(\mathcal{E})=\mu\mathrm{supp}(\mathfrak{A}(\mathcal{E})).$ (7.5)
Proof.
Let $(x,\xi)$ be a point in $T^{*}M$. Consider any $C^{1}$-function with $\phi$ with $d\phi(x)=\xi$. For any $c^{\prime}>0$, the equivariant sheaf $\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{-\phi+c+c^{\prime}>t\geq-\phi+c}$ is almost sent to $R_{c}^{\phi}$ on sufficiently small open subset around $x$ under $\mathfrak{A}$. Hence the test sheaves to estimate the both sides of the desired equality coincide. This completes the proof. ∎
8 Application II: Curved sheaves
In Fukaya category theory, we have to deal with curved complexes, bounding cochain, bulk deformations. In sheaf theory (in the setup of Tamarkin category), introducing such notions is not easy (although we can do them as partly explained in [IK]). Our interpretation of the category $\mathrm{Sh}^{\mathbb{R}}_{\tau>0}(X\times\mathbb{R}_{t})$ as the sheaf category of $\Lambda_{0}$-modules allows us to introduce such deformations easily.
8.1 Curved complex and sheaves
We set $R:=\Lambda_{0}$. Let $V$ be a $\mathbb{Z}$-graded $R$-module and $d$ be a $\deg=1$-endomorphism of $V$. We call such a pair $\mathbb{V}:=(V,d)$, a curved $R$-module complex. We sometimes use the notation $d_{\mathbb{V}}:=d$. We call $d^{2}$ is the curvature of $\mathbb{V}$.
A morphism between a curved complex is a graded $R$-linear morphism between underlying $\mathbb{Z}$-graded $R$-modules. We denote the category of curved $R$-module complexes by $CCh^{c}(R)$.
Let $\mathbb{V}_{i}:=(V_{i},d_{i})$ ($i=1,2$) be curved complexes. The tensor product is defined by the graded tensor product $V_{1}\otimes V_{2}$ equipped with the differential $d_{\mathbb{V}_{1}}\otimes 1+1\otimes d_{\mathbb{V}_{2}}$. This defines a monoidal category structure.
Definition 8.1.
We say a category enriched over $CCh^{c}(R)$ is an $R$-linear curved dg-category.
Let $(V_{i},d_{i})$ be curved complexes. The space of linear maps $\operatorname{Hom}(V_{1},V_{2})$ is again a curved complex $\mathbb{H}om(V_{1},V_{2})$ where its differential is defined by
$f\mapsto d_{V_{2}}\circ f-f\circ d_{V_{1}}.$ (8.1)
Hence $CCh^{c}(R)$ is itself a curved dg-category.
Definition 8.2 (Flat part).
Let $\mathbb{V}$ be a curved complex. Note that $\mathbb{V}_{0}:=(\ker(d^{2}),d)$ is a usual complex. We say $\mathbb{V}_{0}$ is the flat part of $\mathbb{V}$.
Let $(V_{i},d_{i})$ ($i=1,2$) be curved complexes. Considering $\mathbb{H}om(V_{1},V_{2})_{0}$, we obtain the dg-category $CCh(R)$ of curved complexes. This category contains the dg-category of chain complexes of $R$-modules $Ch(R)$ as a subcategory.
Definition 8.3.
Let us consider a sequence of curved complexes:
$\mathbb{V}_{1}\rightarrow\mathbb{V}_{2}\rightarrow\cdots\mathbb{V}_{n}.$ (8.2)
Suppose that this sequence is an exact sequence for each graded part. Such a sequence is called an acyclic complex.
Now we construct the so called totalization complex. For the definition, we refer to [Pos11].
Remark 8.4.
In [Pos11], there are three kinds of derived categories. We choose the one caled “absolute” one. Since the objects we are interested in are finite in some sense, we believe that this choice is not essential for our purpose.
We denote the full subcategory spanned by acyclic complexes by $Acycl$, We set the Drinfeld quotient by
$\mathrm{CCh}(R):=CCh(R)/Acycl.$ (8.3)
Similarly, we consider the category of curved sheaves as a global version of the above story. Let $X$ be a manifold. A curved sheaf $\mathcal{E}$ is a $\mathbb{Z}$-graded sheaf with a degree 1 endomorphism. We consider the category defined by
  1. 1.
    the objects is the curved sheaves
  2. 2.
    a morphism is a graded morphism between graded sheaves underlying curved sheaves.
We denote this category by $CSh^{c}(R_{X})$, which is a curved dg-category. By replacing the hom-spaces by flat parts, we obtain a dg-category $CSh(R_{X})$. We can similarly define the subcategory of the totalization of the acyclic complexes $Acycl$. Then we define
$\mathrm{CSh}(X,R):=CSh(R_{X})/Acycl.$ (8.4)
Associated to $\mathrm{CSh}(X,R)$, the derived dg category of curved sheaves.
8.2 Twisted sheaves
Inside $\mathrm{CSh}(X,R)$, there exists a well-behaved subclass of objects: weakly unobstructed sheaves. For any object $\mathcal{E}\in\mathrm{CSh}(X,R)$, we have $\mathcal{E}\otimes R_{X}\cong\mathcal{E}$ under derived tensor product. This implies that we have a morphism
$w\colon\mathcal{E}\rightarrow\mathcal{E}[2]$ (8.5)
associated to each $w\in H^{2}(X,R)$.
Definition 8.5.
Take $w\in H^{2}(X,R)$. We denote the full subcategory of $\mathrm{CSh}(X,R)$ consisting of the objects whose curvature is cohomologically $w$ by $\mathrm{Sh}(X,R,w)$.
We first note that we have the following isomorphism:
[…]
Recall that $\mathfrak{m}$ is the maximal ideal of $R$. Take a cohomology $e^{w}\in H^{2}(X,1+\mathfrak{m})$. Fix a Cech 2-cocycle $e^{c_{ijk}}\in 1+\mathfrak{m}$. We consider the category consisting of objects as follows:
  1. 1.
    For each $U_{i}$, we have an object $\mathcal{E}_{i}$ of $\mathrm{Sh}(U_{i},R)$,
  2. 2.
    On the restriction to $U_{i}\cap U_{j}$, we have a specified isomorphism $\mathcal{E}_{i}\cong\mathcal{E}_{j}$ in $\mathrm{Sh}(U_{i}\cap U_{j},R)$.
  3. 3.
    On the restriction to $U_{i}\cap U_{j}\cap U_{k}$, the associated automorphism of $\mathcal{E}_{i}$ is $e^{c_{ijk}}$.
The resulting category does not depend on the choice of Cech representative of an element $e^{w}\in H^{2}(X,1+\mathfrak{m})$. We denote the resulting category by $\mathrm{Sh}_{tw}(X,R,e^{w})$.
We now deduce that the above two categories are just two presentations of the same category.
Theorem 8.6.
For $w\in H^{2}(X,\mathfrak{m})$, we have $\mathrm{Sh}_{tw}(X,R,e^{w})\simeq\mathrm{Sh}(X,R,w)$.
Proof.
For the sake of simplicity, we assume $\mathbb{K}=\mathbb{C}$.
Take a good cover $\{U_{i}\}$. Over each $U_{i}$, the restriction $w|_{U_{i}}$ has a primitive 1-form $\alpha_{i}$.
Take an object $\mathcal{E}\in\mathrm{Sh}(X,R,w)$. Twisting by $\alpha_{i}$ gives an equivalence of $\mathcal{E}|_{U_{i}}\in\mathrm{Sh}(U_{i},R,w|_{U_{i}})\cong\mathrm{Sh}(U_{i},R)\ni\mathcal{E}_{i}$. On the overlap $U_{i}\cap U_{j}$, we have an isomorphism given by
$\mathcal{E}_{i}\xrightarrow{\times e^{f_{ij}}}\mathcal{E}_{j}$ (8.7)
where $f_{ij}\in C^{\infty}(U_{i}\cap U_{j},\Lambda_{+})$ is a primitive of $\alpha_{i}-\alpha_{j}$. As in the usual Cech–de Rham isomorphism, the composition $e^{f_{ij}}e^{f_{jk}}e^{f_{ki}}$ is a constant and given by $e^{c_{ijk}}$ where $c_{ijk}$ is a Cech representative of $w$. This completes the proof.
For general coefficients, the same proof wokrs by replacing the de Rham resolution with Cech resolution. ∎
Example 8.7 (Curved connection).
There exists a category closely related to the above idea. Note that, over the field $\mathbb{C}$, the isomorphism $e^{(-)}$ is extended to
$e^{(-)}\colon\Lambda_{0}\rightarrow\mathbb{C}^{*}+\mathfrak{m}$ (8.8)
where the RHS is the units of $\Lambda_{0}$.
Let $\mathcal{E}$ be a $C^{\infty}$-module with a flat connection. If $\mathcal{E}$ is associated to a vector bundle, it is well-known that the flat sections form a locally constant sheaf and the assignment gives an equivalence.
Similarly, if $\mathcal{E}$ is $C^{\infty}$-module with a connection whose curvature is $w\in\Omega^{2}(X,R)$, they form a dg-category. Then, the above theorem tells us that the category of vector bundles with connections such that the curvature $w$ can be embedded into the category of twisted sheaves.
This construction should be related to the B-field deformation/bulk deformation of Fukaya category.
8.3 Twisted sheaf quantization and bounding cochain
In this section, we explain how we can run the theory in [IK] in the twisted setup. For the details, we refer to [IK].
Note that any object in $\mathrm{Sh}(X,R,w)$ can be locally viewed as an object of $\mathrm{Sh}(X,R)$, one can define $\mu\mathrm{supp}$. Similarly, we say an object $\mathcal{E}$ in $\mathrm{Sh}(X,R,w)$ is a sheaf quantization if it is locally a sheaf quantization viewed as an object of $\mathrm{Sh}(X,R)$.
Also, the low-energy standard sheaf quantization construction has local nature, we get a low-energy sheaf quantization in $\mathrm{Sh}(X,R,w)$ for any Lagrangian brane. Then, in the exactly the same way, one can construct a curved dga associated to a Lagrangian brane. An existence of a Maurer–Cartan element implies an existence of sheaf quantization in $\mathrm{Sh}(X,R,w)$.
Remark 8.8.
It is even possible to formulate curved sheaf quantizations: namely, instead of considering curved twisted complex of hom-spaces as in [IK], one can directly construct a curved twisted complex of sheaves in $\mathrm{CSh}(X,R)$. Then one can formulate the Maurer–Cartan equation in [IK] as a Maurer–Cartan equation of the curved sheaf itself. The resulting theory is obviously the same one obtained in [IK].